OpenStudy (anonymous):

Find the slope of the function f(x)= sqrt(5x) at x=5

OpenStudy (anonymous):

\[f(x) = \sqrt{5x}\]

OpenStudy (anonymous):

You have to find the derivative of f(x). Then you have to calculate the value of the derivative for x=5, that is the f'(5).

OpenStudy (anonymous):

How do I do that? Like how do I find the derivative?

OpenStudy (anonymous):

Would I substitute in 5? That would give me f(x) = 5

OpenStudy (anonymous):

Have you learned derivatives?

OpenStudy (anonymous):

Yes I have. but I dont know what I'm doing, to be honest. I'm so lost in class. But Wouldn't you have to put sqrt5x into 5? thats what your supposed to do, right? but you can't. thats why I'm confused

OpenStudy (anonymous):

To find the slope, you have to find the derivative of the function first. The derivative is denoted by f'(x). For \[f(x) = \sqrt{5x} = \sqrt{5}\sqrt{x} \] you have: \[f'(x) = (f(x))' = (\sqrt{5}\sqrt{x})' = \sqrt{5}(\sqrt{x})' = \sqrt{5} \frac{1}{2\sqrt{x}} = \frac{\sqrt{5}}{2\sqrt{x}}\] Then to find the slope at x=5, you have to find f'(5), that is, in the derivative formula you replace x with 5.

OpenStudy (anonymous):

so would I end up with 25/2

OpenStudy (anonymous):

\[f'(5) = \frac{\sqrt{5}}{2\sqrt{5}} = ..\]

OpenStudy (anonymous):

whoops, that would make it 1/2

OpenStudy (phi):

are you doing multiple guess ? you could write \[ \sqrt{5x} = \sqrt{5} x^{\frac{1}{2}} \] and use the rule \[ \frac{ d x^n}{dx} = n x^{n-1} \]

OpenStudy (anonymous):

That's right @cutie.patootie !

OpenStudy (anonymous):

Oh sweet! Thanks!!!

OpenStudy (anonymous):

that was so confusing.. Lol

OpenStudy (anonymous):

Do you know what should be done if you want to find the slope at x=20 now?

OpenStudy (anonymous):

I would just replace x with 20, right?

OpenStudy (anonymous):

Yes, you would replace x with 20 in the f'(x) formula. So what would you get?

OpenStudy (anonymous):

1/10

OpenStudy (anonymous):

\[f'(x) = \frac{\sqrt{5}}{2\sqrt{x}} \] Replacing x with 20: \[f'(20) = \frac{\sqrt{5}}{2\sqrt{20}} \] \[f'(20) = \frac{\sqrt{5}}{2\sqrt{4*5}} \] \[f'(20) = \frac{\sqrt{5}}{2\sqrt{4}*\sqrt{5}} \] \[f'(20) = \frac{\sqrt{5}}{2*2\sqrt{5}} \] \[f'(20) = \frac{\sqrt{5}}{4\sqrt{5}} \] \[f'(20) = \frac{1}{4}\]

OpenStudy (anonymous):

Ohhh. I didn't separate the 20

OpenStudy (anonymous):

It's ok, you will get used to it. In general you want to siplify the square root, so you think how you can separate the value inside the square root, so you can pull something out of the root. The square root of 4 is 2, so you could write 20=4*5.

OpenStudy (anonymous):

okay, thank you for your help:))))

OpenStudy (anonymous):

Anytime!